The Alpine House, Inc, is a large retailer of snow skis. The company assembled the information shown below for the quarter ended March 31: Amount Sales Selling price per pair of skis Variable selling expense per pair of skis Variable administrative expense per pair of skis Total fixed selling expense Total fixed administrative expense Beginning merchandise inventory Ending merchandise inventory Merchandise purchases $ 1,144,000 440 47 18 150,880 s110,e00 S 78,800 $ 1e5,eee s300,880 Required 1. Prepare a traditional Income statement for the quarter ended March 31 2. Prepare a contribution format income statement for the quarter ended March 31 3. What was the contribution margin per unit? Complete this question by entering your answers in the tabs below Required 1 Required 2 Required 3 Prepare a traditional income statement for the quarter ended March 31. The Alpine House, Inc. Traditional Income Statement Selling and administrative expensers
rchandise purchases 30e, eee Required 1. Prepare a traditional income statement for the quarter ended March 31 2 Prepare a contribution format income statement for the quarter ended March 31. 3. What was the contribution margin per unit? Complete this question by entering your answers in the tabs below. Required 1Required 2 Required 3 Prepare a contribution format income statement for the quarter ended March 31. The Alpine House, Inc. Contribution Format Income Statement Variable expenses: Fixed expenses Required 1 Required 3 > arch
Required 1 Required 2 Required 3 what was the contribution margin per unit? (Round your final answer to nearest whole dolilar.) Required 2
Answer #1

Answers

  • All working forms part of the answer
  • Working for Requirement 1 and Requirement 2

A

Beginning Merchandising Inventory

$                70,000

B

Merchandise Purchases

$              300,000

C

Ending merchandise Inventory

$              105,000

D = A + B - C

Cost of Goods Sold

$              265,000

Sales = $ 1144000, Sale price = $ 440 per unit
Hence, units sold = 1144000 / 440 = 2,600 units

  • Requirement 1

--Traditional Income Statement

Sales

$         1,144,000

Cost of Goods Sold

$             265,000

Gross Profits

$             879,000

Selling & administrative expenses:

Variable selling expense [2600 x x$ 47]

$              122,200

Variable administrative expense [2600 x x$ 18]

$                46,800

Total Fixed selling expense

$              150,000

Total Fixed administrative expense

$              110,000

$             429,000

Net Operating Income

$             450,000

  • Requirement 2

--Contribution format Income Statement

Sales

$         1,144,000

Variable expenses:

Cost of Goods Sold

$              265,000

Variable selling expense [2600 x x$ 47]

$              122,200

Variable administrative expense [2600 x x$ 18]

$                46,800

$             434,000

Contribution margin

$             710,000

Fixed expenses:

Total Fixed selling expense

$              150,000

Total Fixed administrative expense

$              110,000

$             260,000

Net Operating Income

$             450,000

  • Requirement 3

Contribution margin = $ 710,000 [see Requirement 2]
Units sold = 2600 units [See Working]
Contribution margin per unit = 710000 / 2600 = $ 273.08

Answer = $ 273